OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -1

OpenStudy (jack1):

rearrange 2x + y = 1 into y = 1 - 2x then sub into 4x + 2y = -1

OpenStudy (jack1):

4x + 2y = -1 4x + 2(1 - 2x) = -1 can you solve from here...?

OpenStudy (anonymous):

yes thank you